# Self-gravitating Higgs field of scalar charge

### 26 Dec 2023 v1

@@ Abstract
The self-gravitating Higgs field of a scalar charge has been studied. It is shown that in the zero and first approximation of the smallness of the scalar charge, the gravitational field of the scalar charge is described by the Schwarzschild-de Sitter metric with a cosmological constant determined by the vacuum potential of the Higgs field. An equation for the perturbation of the vacuum potential is obtained and studied. Particular exact solutions of the field equation are given. It is shown that in the case of a naked singularity, solutions to the field equation have the character of microscopic oscillations with a Compton wavelength. Limiting asymptotic cases of the behavior of solutions are studied and their comparative analysis is carried out in relation to the Fisher solution. The averaging of microscopic oscillations of the scalar field was carried out and it was shown that they make a negative contribution to the macroscopic energy of the scalar field, reducing the observed value of the Black Hole mass. A computer simulation of a scalar field has been carried out, demonstrating various types of behavior of solutions.
@@ 4.1 Field equations
$2r\Phi^{\prime 2}+(\lambda+\nu)^{\prime}=0;$ (31)
@@
$\displaystyle\frac{d^{2}\phi}{dr^{2}}+\frac{d}{dr}\ln\bigl{(}r^{2}\mathrm{e}^{\nu_{0}(r)}\bigr{)}\ frac{d\phi}{dr}+2m^{2}_{s}\phi=0.$ (37)
@@ 4.2 Horizons and singularity
$\begin{array}[]{llll}\gamma\leqslant 0:&\mathbf{X_{1}}=[0,x_{1}],&\mathbf{X_{3}}=(x_{1},+\infty),&\\[12.0pt] \displaystyle 0<\gamma<\frac{4}{27}:&\mathbf{X_{1}}=[0,x_{1}],&\mathbf{X_{2}}=(x_{1},x_{2}),&\mathbf{X_{3}}=(x_{2},+\infty),\\[12.0pt] \displaystyle\gamma>\frac{4}{27}:&\mathbf{X_{3}}=[0,+\infty).&&\\ \end{array}$ (49)
@@ 6 Averaging of scalar potential oscillations
Taking into account the rapidly oscillating nature of the solutions to the system of equations (40) for $\sigma x\gg 1$, let us average the values (74) over a sufficiently large interval of the radial variable, using the technique of averaging cosmological fluctuations of the scalar field [13] - - [14]. Namely, let us introduce the macroscopic average of the rapidly varying function $f(r)$:
@@
$\left.\overline{\varepsilon}\ \right|_{\gamma x^{3}\to\infty}\backsimeq-\frac{\gamma\sigma^{2}|\phi_{0}|^{2}}{16\pi(2m)^{2}x^{2}};\qquad\overline{\varepsilon}+\overline{p_{\parallel}}\to 0.$
@@
increases in proportion to the radius, thereby reducing the observed mass of the Black Hole.
@@ 7.2 Small values of the parameter $\gamma<4/27$
$\mathbf{X_{1}}:=[0,r_{1});\quad\mathbf{X_{2}}=(x_{1},x_{2});\quad\mathbf{X_{3}}=(x_{2},+\infty).$
@@ 7.3 Large values of the parameter $\gamma>4/27$
At $\gamma>4/27$ there are no horizons in the (36) metric, i.e., we have a Black Hole with a bare singularity $r=0$. In this case, the field equations (40) have analytical solutions in the entire space $r\geqslant 0$. Figure 7.3 and LABEL:Ignatev5 show graphs of the potential function $\psi(x)$ and its derivative $\psi^{\prime}(x)$ at
@@ Conclusion
  • The macroscopic average energy densities and pressures of scalar oscillations are calculated and it is shown that the total macroscopic energy density of oscillations is negative, and the total equation of state corresponds to macroscopic phantom scalar field.
  • It is shown that the negative macroscopic oscillation energy density leads to a decrease in the observed mass of the Black Hole.

### 05 Mar 2024 v2

@@ Abstract
The self-gravitating Higgs field of a scalar charge has been studied. It is shown that in the zero and first approximation of the smallness of the scalar charge, the gravitational field of the scalar charge is described by the Schwarzschild-de Sitter metric with a cosmological constant determined by the vacuum potential of the Higgs field. An equation for the perturbation of the vacuum potential is obtained and studied. Particular exact solutions of the field equation are given. It is shown that in the case of a naked singularity, solutions to the field equation have the character of microscopic oscillations with a Compton wavelength. Limiting asymptotic cases of the behavior of solutions are studied and their comparative analysis is carried out in relation to the Fisher solution. The averaging of microscopic oscillations of the scalar field was carried out and it was shown that at $\Lambda>0$ they make a negative contribution to the macroscopic energy of the scalar field, reducing the observed value of the Black Hole mass. A computer simulation of a scalar field has been carried out, demonstrating various types of behavior of solutions.
@@ 4.1 Field equations
$r\Phi^{\prime 2}+(\lambda+\nu)^{\prime}=0;$ (31)
@@
$\displaystyle\frac{d^{2}\phi}{dr^{2}}+\frac{d}{dr}\ln\bigl{(}r^{2}\mathrm{e}^{\nu_{0}(r)}\bigr{)}\frac{d\phi}{dr}+2m^{2}_{s}\phi=0.$ (37)
@@ 4.2 Horizons and singularity
$\begin{array}[]{llll}\gamma\leqslant 0:&\mathbf{X_{1}}=[0,x_{1}],(\mathrm{T});&\mathbf{X_{3}}=(x_{1},+\infty),(\mathrm{R});&\\[12.0pt] \displaystyle 0<\gamma<\frac{4}{27}:&\mathbf{X_{1}}=[0,x_{1}],(\mathrm{T});&\mathbf{X_{2}}=(x_{1},x_{2}),(\mathrm{R});&\mathbf{X_{3}}=(x_{2},+\infty),(\mathrm{T});\\[12.0pt] \displaystyle\gamma>\frac{4}{27}:&\mathbf{X_{3}}=[0,+\infty),(\mathrm{T}).&&\\ \end{array}$ (49)
@@ 6 Averaging of scalar potential oscillations
Taking into account the rapidly oscillating nature of the solutions to the system of equations (40) for $\sigma x\gg 1$, let us average the values (74) over a sufficiently large interval of the radial variable, using the technique of averaging cosmological fluctuations of the scalar field [12][13]. Namely, let us introduce the macroscopic average of the rapidly varying function $f(r)$:
@@
$\left.\overline{\varepsilon}\ \right|_{\gamma x^{3}\to\infty}\backsimeq-\frac{\gamma\sigma^{2}|\phi_{0}|^{2}}{16\pi(2m)^{2}x^{2}};\qquad\overline{p_{\parallel}}\to\overline{\varepsilon}.$
@@
increases in proportion to the radius, thereby reducing the observed mass of the Black Hole. Note, however, that according to (49) the $\forall\gamma$ region (see (72)) for $\gamma>0$ is a $\mathrm{T}$-region, in for which there is no stationary state and there is no infinitely distant observer. In fact, in $\mathrm{T}$-regions we must swap spatial and temporal coordinates. Therefore, for $\gamma>0$, the <<external region>> $\forall\gamma$ is, in fact, a cosmological continuation of the solution. Thus, to correctly interpret the static scalar charge field, we must set $\gamma<0$. In this case, a possible and correct interpretation of the formulas (6) and those following them is that the macroscopic oscillation energy density becomes positive, and at the same time the macroscopic mass of the scalar halo also becomes positive. We intend to return to a more detailed study of this issue in the next part of the article.
@@ 7.2 Small values of the parameter $\gamma<4/27$
$\mathbf{X_{1}}=[0,r_{1});\quad\mathbf{X_{2}}=(x_{1},x_{2});\quad\mathbf{X_{3}}=(x_{2},+\infty).$
@@ 7.3 Large values of the parameter $\gamma>4/27$
At $\gamma>4/27$ there are no horizons in the (36) metric, i.e., we have a Black Hole with a bare singularity $r=0$. In this case, the field equations (40) have analytical solutions in the entire space $r\geqslant 0$. Figure 7.3 and 7.3 show graphs of the potential function $\psi(x)$ and its derivative $\psi^{\prime}(x)$ at
@@ Conclusion
  • The macroscopic average energy densities and pressures of scalar oscillations are calculated and it is shown that the total macroscopic oscillation energy density is negative for $\Lambda>0$ and positive for $\Lambda<0$.
  • It is shown that the macroscopic oscillation energy density leads to a change in the observed mass of the black hole.